# FlowTok: Flowing Seamlessly Across Text and Image Tokens

### v2

@@ 5.1 Implementation Details
Image Tokenizer. We build our image tokenizer upon the official TA-TiTok [40] codebase with minimal modifications. The encoder uses ViT-B [23], while the decoder uses ViT-L, both operating with a patch size of . To align with the output sequence length of CLIP’s text encoder, we set the number of 1D latent tokens to 77 and the token dimension to 16. Additionally, we enhance the tokenizer with RoPE [76] and SwiGLU FFN [71]. Notably, our enhanced tokenizer achieves a FID of 1.02 in a zero-shot evaluation on the ImageNet validation set, matching the performance of the original TA-TiTok with 128 tokens.
Text Projector. We train a text projector for text-to-image generation, which transforms CLIP text embeddings into a latent representation of shape , aligning with the image latent space encoded by our image tokenizer. The text projector consists of six Transformer [81] blocks, each comprising a multi-head self-attention mechanism and a multi-layer perceptron (MLP), both enhanced with skip connections [34] to ensure stable training.
Text Decoder. We train a text decoder for image-to-text generation, composed of six Transformer [81] blocks, similar to the text projector. The decoder takes the text latent representation as input and outputs the corresponding CLIP text tokenizer indices, which can be further converted into text using the CLIP text tokenizer.
FlowTok. We adopt DiT [61] blocks as the fundamental building units of our FlowTok to model token interactions. Specifically, we follow the DiT architecture to implement FlowTok-B for efficient ablation studies and FlowTok-XL for enhanced performance. To further push the performance, we scale up the depth, width, and number of attention heads, constructing FlowTok-H with 1.1B parameters. The detailed model configurations are provided in Tab. 1.
@@
| GLIDE [58] | 5.0B | ✗ | - | - | 12.24 | - |
| Dalle2 [66] | 6.5B | ✗ | - | - | 10.39 | - |
| LlamaGen [78] | 775M | ✗ | - | - | - | 25.59 |
| PixArt- [16] | 630M | ✗ | 94.1 | 7.9 | 7.32 | 9.85 |
| SDXL [62] | 2.6B | ✗ | - | - | - | 8.76 |
| LDM [69] | 1.4B | ✓ | - | - | 12.63 | - |
| Stable-Diffusion-1.5 [69] | 860M | ✓ | 781.2 | - | 9.62 | - |
| Stable-Diffusion-2.1 [69] | 860M | ✓ | 1041.6 | - | 13.45 | 26.96 |
| Show-o [85] | 1.3B | ✓ | - | 1.0 | 9.24 | 14.99 |
@@
Dataset. We employ open-source datasets [40] to facilitate the reproducibility of FlowTok’s simple framework. Specifically, our image tokenizer is trained on DataComp-1B [29], and our text tokenizer is trained on COCO [49]. For text-to-image generation, inspired by recent works [21, 16, 15, 93], we adopt a two-stage training strategy: pre-training and fine-tuning. The pre-training stage leverages a combination of DataComp-1B [29], CC12M [14], and LAION-aesthetic [1], while the fine-tuning stage incorporates additional high-quality datasets, including LAION-art [3], LAION-pop [4], JourneyDB [77], and DALLE3-1M [25]. For image-to-text generation, we follow the Karpathy split [39] of COCO [19] to divide the training and validation sets. Detailed dataset information is provided in the Appendix.
@@
Evaluation. We follow standard evaluation practices to report relevant metrics for both text-to-image and image-to-text generation. Specifically, for text-to-image generation, we report FID-30K on the COCO [49], and FID on MJHQ-30K [45]. For image-to-text generation, we report BLEU-4 [60], METEOR [10], ROUGE [48], CIDEr [82], and SPICE [8] on the COCO Karpathy Split [39]. To incorporate classifier-free guidance (CFG) [35] within FlowTok, we follow CrossFlow[52] and utilize a CFG indicator. Unless otherwise stated, we find that using only 20 steps for sampling is sufficient due to the small 1D latent shape of FlowTok. This significantly speeds up the inference process, enabling faster generation without compromising performance.
@@
| MNIC [31] | 30.9 | 27.5 | 55.6 | 108.1 | 21.0 |
| MIR [43] | 32.5 | 27.2 | - | 109.5 | 20.6 |
| NAIC-CMAL [33] | 35.3 | 27.3 | 56.9 | 115.5 | 20.8 |
| SATIC [95] | 32.9 | 27.0 | - | 111.0 | 20.5 |
| SCD-Net [56] | 37.3 | 28.1 | 58.0 | 118.0 | 21.6 |
@@ 5.2 Main Results
Text-to-Image Generation. We report zero-shot text-to-image generation results on COCO [49] and MJHQ-30K [45] in Tab. 2. The compared methods are categorized into two groups: text as conditions, where text serves as a guiding signal for image generation, and text as source distributions, where text is directly modeled as a distribution in the generative process. As observed, FlowTok achieves comparable performance to prior methods in both categories on COCO FID-30K. Specifically, compared to CrossFlow [52], which also uses text as the source distribution, FlowTok-H attains a FID-30K of 9.67—roughly on par with CrossFlow. When further evaluating FlowTok on MJHQ-30K to assess the aesthetic quality of generated images, we find that, despite being trained solely on publicly available datasets without access to high-quality proprietary data, FlowTok-XL already surpasses other state-of-the-art models, demonstrating its ability to generate diverse, high-quality images. Furthermore, FlowTok-H further improves the FID score to 7.15, underscoring its superior image generation capabilities.
Beyond performance, FlowTok requires significantly fewer training resources compared to existing state-of-the-art models. Specifically, FlowTok-XL completes training in just 20.4 8-A100 days, while FlowTok-H increases the budget slightly to 26.1 8-A100 days. In contrast, the most efficient text-as-condition model, PixArt- [16], still demands 94.1 8-A100 days. Compared to CrossFlow [52], which also treats text as source distributions and requires 78.8 8-A100 days, FlowTok is much more efficient.
@@
Image-to-Text Generation. We evaluate image-to-text generation on COCO [49] using the Karpathy split [39], with results summarized in Tab. 3. To ensure a fair comparison, we categorize methods into two groups: direct flow from image to text distributions, which represents a new paradigm leveraging flow matching for direct image-to-text transformation, and other methods, considering only those trained without CIDEr optimization. Within the direct flow category, FlowTok-XL consistently outperforms its counterpart, CrossFlow [52], across most metrics. Specifically, FlowTok-XL achieves a BLEU-4 (B@4) score of 37.1, surpassing CrossFlow by 0.7, and a CIDEr score of 117.0, exceeding CrossFlow by 0.8. Moreover, compared to state-of-the-art methods from other paradigms, FlowTok-XL demonstrates competitive performance, highlighting direct flow matching as a promising approach for image-to-text generation. Notably, FlowTok performs image-to-text generation under the same formulation using compact 1D tokens, theoretically requiring fewer training resources and enabling faster sampling compared to paradigms that operate on 2D latents, as adopted by CrossFlow. However, a direct quantitative comparison is not possible, as CrossFlow has not released the corresponding checkpoint for evaluation.
@@ 5.3 Ablation Studies
Text Alignment Target. We first investigate the choice of alignment target for the projected text tokens in Tab. LABEL:tab:ablation_loss_target. A straightforward baseline is to directly apply average pooling (row 1) to the original CLIP text embedding along the channel dimension, reducing the dimensionality from 768 to 16 to match . However, this approach performs significantly worse compared to using a simple MLP to learn the alignment target (row 2), as inspired by prior works [32, 59, 92]. We attribute this performance gap to the fact that adjacent channels in the CLIP text embedding are not necessarily correlated, and simple average pooling discards too much semantic information. In contrast, a learnable MLP mitigates this information loss, making it a more effective choice for defining the text alignment target.
Text Alignment Loss Function. Next, we examine the text alignment loss function in Tab. LABEL:tab:ablation_loss_function. Besides the contrastive loss adopted from [65], we explore using a cosine similarity loss, similar to [87]. Specifically, we compute the cosine similarity between the text tokens and the alignment target, applying a penalty to pairs with similarity below a threshold. Our experiments show that while both loss functions are effective, the contrastive loss achieves better performance.
@@ 6 Conclusion
Appendix
The supplementary material provides additional information:
A More Implementation Details
| model | dataset | filtering | ||
| resolution | aesthetic | watermark | ||
| Text Decoder | COCO [49] | |||
| Image Tokenizer | DataComp [29] | ✓ | ||
| FlowTok: pre-training | DataComp [29] | ✓ | ✓ (5.0) | ✓ |
| CC12M [14] | ✓ | ✓ (5.0) | ✓ | |
| LAION-aesthetic [1] | ✓ | ✓ | ||
| FlowTok: fine-tuning | [29] | ✓ | ✓ (6.0) | ✓ |
| [3] | ✓ | ✓ | ||
| [4] | ✓ | ✓ | ||
| DALLE3-1M [25] | ||||
| JourneyDB [77] | ||||
Dataset Filtering. In line with previous works [40], we apply three filtering criteria to curate open data for training the image tokenizer and FlowTok: resolution, aesthetic quality, and watermark filtering. The COCO [49, 19] dataset is used directly to train the text decoder without any filtering. Details of the applied filtering criteria are shown in Tab. 5.
Specifically, resolution filtering is applied during the training of the image tokenizer and for text-to-image generation. This ensures that the longer side of each image is at least 256 pixels and the aspect ratio is below 2. For text-to-image training, we further apply aesthetic filtering using the LAION-aesthetic-v2 predictor [2] to retain only high-quality images. Images with aesthetic scores above 5.0 are retained during the pre-training stage, while a stricter threshold of 6.0 is used during fine-tuning to ensure even higher image quality.
Additionally, watermark filtering is implemented for FlowTok’s text-to-image generation by using the LAION-WatermarkDetector [5], removing images with watermark probabilities exceeding 0.5. Synthetic datasets such as JourneyDB [77] and DALLE3-1M [25] are exempt from these filtering steps, as they inherently meet our high resolution and quality standards.
| hyper-parmeters | pre-training | fine-tuning |
| optimizer | AdamW | AdamW |
| optimizer- | 0.9 | 0.9 |
| optimizer- | 0.95 | 0.95 |
| weight decay | 0.03 | 0.03 |
| lr | 0.0004 | 0.0002 |
| lr scheduling | constant | constant |
| lr warmup steps | 10K | 0 |
| batch size | 4096 | 4096 |
| training steps | 250K | 150k |
Training Hyper-parameters. Tab. 6 provides the complete list of hyper-parameters used for training FlowTok.
B Qualitative Examples of FlowTok
Additional Generation Results. Fig. 5, Fig. 6, and Fig. 7 present additional text-to-image generation results produced by FlowTok, demonstrating its ability to generate diverse, high-fidelity images. Meanwhile, Fig. 8 displays the image-to-text generation results, showcasing FlowTok’s capability to produce accurate and descriptive captions.
C Limitations and Future Work
The primary limitation of FlowTok arises during text-to-image generation. To match the compact dimensionality of image latents (e.g., 16), FlowTok projects CLIP text embeddings into the same low-dimensional latent space. While the text alignment loss helps preserve semantic information, some degree of information loss is inevitable during this projection. Consequently, the alignment between text and generated images may be weaker compared to state-of-the-art models employing cross-attention mechanisms. To address this, one potential solution is to introduce a stronger alignment loss that better retains textual semantics. A more fundamental approach, however, involves increasing the channel dimensionality of image latents by aligning them with vision foundation models [59] during image tokenizer training, as suggested by VA-VAE [88]. This strategy aims to identify an optimal channel dimension for both image and text tokens, achieving a balance between preserving semantic information and maintaining efficiency in training and inference.
Additionally, FlowTok currently utilizes only the vanilla flow matching technique to validate the framework’s effectiveness. However, many recent advancements in flow matching, such as logit-normal sampling [26], have not yet been explored in our model. Incorporating these techniques could accelerate convergence and enhance performance.
Finally, FlowTok serves as a starting point for exploring efficient direct evolution between text and image modalities. In the future, we aim to extend to a more general framework that can accommodate a broader range of modalities, supporting additional tasks under the same unified formulation.

### v3

@@ 5.1 Implementation Details
Image Tokenizer. We build our image tokenizer upon the official TA-TiTok [kim2025democratizing] codebase with minimal modifications. The encoder uses ViT-B [dosovitskiy2020image], while the decoder uses ViT-L, both operating with a patch size of . To align with the output sequence length of CLIP’s text encoder, we set the number of 1D latent tokens to 77 and the token dimension to 16. Additionally, we enhance the tokenizer with RoPE [su2021roformer] and SwiGLU FFN [shazeer2020glu]. Notably, our enhanced tokenizer achieves a FID of 1.02 in a zero-shot evaluation on the ImageNet validation set, matching the performance of the original TA-TiTok with 128 tokens.
Text Projector. We train a text projector for text-to-image generation, which transforms CLIP text embeddings into a latent representation of shape , aligning with the image latent space encoded by our image tokenizer. The text projector consists of six Transformer [vaswani2017attention] blocks, each comprising a multi-head self-attention mechanism and a multi-layer perceptron (MLP), both enhanced with skip connections [he2016deep] to ensure stable training.
Text Decoder. We train a text decoder for image-to-text generation, composed of six Transformer [vaswani2017attention] blocks, similar to the text projector. The decoder takes the text latent representation as input and outputs the corresponding CLIP text tokenizer indices, which can be further converted into text using the CLIP text tokenizer.
FlowTok. We adopt DiT [peebles2023scalable] blocks as the fundamental building units of our FlowTok to model token interactions. Specifically, we follow the DiT architecture to implement FlowTok-B for efficient ablation studies and FlowTok-XL for enhanced performance. To further push the performance, we scale up the depth, width, and number of attention heads, constructing FlowTok-H with 1.1B parameters. The detailed model configurations are provided in Tab. 1.
@@
| GLIDE [nichol2021glide] | 5.0B | ✗ | - | - | 12.24 | - |
| Dalle2 [ramesh2022hierarchical] | 6.5B | ✗ | - | - | 10.39 | - |
| LlamaGen [sun2024autoregressive] | 775M | ✗ | - | - | - | 25.59 |
| PixArt- [chen2024pixartalpha] | 630M | ✗ | 94.1 | 7.9 | 7.32 | 9.85 |
| SDXL [podell2023sdxl] | 2.6B | ✗ | - | - | - | 8.76 |
| LDM [rombach2022high] | 1.4B | ✓ | - | - | 12.63 | - |
| Stable-Diffusion-1.5 [rombach2022high] | 860M | ✓ | 781.2 | - | 9.62 | - |
| Stable-Diffusion-2.1 [rombach2022high] | 860M | ✓ | 1041.6 | - | 13.45 | 26.96 |
| Show-o [xie2024show] | 1.3B | ✓ | - | 1.0 | 9.24 | 14.99 |
@@
Dataset. We employ open-source datasets [kim2025democratizing] to facilitate the reproducibility of FlowTok’s simple framework. Specifically, our image tokenizer is trained on DataComp-1B [gadre2024datacomp], and our text tokenizer is trained on COCO [lin2014microsoft]. For text-to-image generation, inspired by recent works [dai2023emuenhancingimagegeneration, chen2024pixartalpha, chen2024pixart, zheng2024cogview3finerfastertexttoimage], we adopt a two-stage training strategy: pre-training and fine-tuning. The pre-training stage leverages a combination of DataComp-1B [gadre2024datacomp], CC12M [changpinyo2021conceptual], and LAION-aesthetic [laion-aesthetic], while the fine-tuning stage incorporates additional high-quality datasets, including LAION-art [laion-art], LAION-pop [laion-pop], JourneyDB [sun2024journeydb], and DALLE3-1M [Egan_Dalle3_1_Million_2024]. For image-to-text generation, we follow the Karpathy split [karpathy2015deep] of COCO [chen2015microsoft] to divide the training and validation sets. Detailed dataset information is provided in the Appendix.
@@
Evaluation. We follow standard evaluation practices to report relevant metrics for both text-to-image and image-to-text generation. Specifically, for text-to-image generation, we report FID-30K on the COCO [lin2014microsoft], and FID on MJHQ-30K [li2024playground]. For image-to-text generation, we report BLEU-4 [papineni2002bleu], METEOR [banerjee2005meteor], ROUGE [lin2004rouge], CIDEr [vedantam2015cider], and SPICE [anderson2016spice] on the COCO Karpathy Split [karpathy2015deep]. To incorporate classifier-free guidance (CFG) [ho2022classifier] within FlowTok, we follow CrossFlow[liu2024flowing] and utilize a CFG indicator. Unless otherwise stated, we find that using only 20 steps for sampling is sufficient due to the small 1D latent shape of FlowTok. This significantly speeds up the inference process, enabling faster generation without compromising performance.
@@
| MNIC [gao2019masked] | 30.9 | 27.5 | 55.6 | 108.1 | 21.0 |
| MIR [lee2018deterministic] | 32.5 | 27.2 | - | 109.5 | 20.6 |
| NAIC-CMAL [guo2020non] | 35.3 | 27.3 | 56.9 | 115.5 | 20.8 |
| SATIC [zhou2021semi] | 32.9 | 27.0 | - | 111.0 | 20.5 |
| SCD-Net [luo2023semantic] | 37.3 | 28.1 | 58.0 | 118.0 | 21.6 |
@@ 5.2 Main Results
Text-to-Image Generation. We report zero-shot text-to-image generation results on COCO [lin2014microsoft] and MJHQ-30K [li2024playground] in Tab. 2. The compared methods are categorized into two groups: text as conditions, where text serves as a guiding signal for image generation, and text as source distributions, where text is directly modeled as a distribution in the generative process. As observed, FlowTok achieves comparable performance to prior methods in both categories on COCO FID-30K. Specifically, compared to CrossFlow [liu2024flowing], which also uses text as the source distribution, FlowTok-H attains a FID-30K of 9.67—roughly on par with CrossFlow. When further evaluating FlowTok on MJHQ-30K to assess the aesthetic quality of generated images, we find that, despite being trained solely on publicly available datasets without access to high-quality proprietary data, FlowTok-XL already surpasses other state-of-the-art models, demonstrating its ability to generate diverse, high-quality images. Furthermore, FlowTok-H further improves the FID score to 7.15, underscoring its superior image generation capabilities.
Beyond performance, FlowTok requires significantly fewer training resources compared to existing state-of-the-art models. Specifically, FlowTok-XL completes training in just 20.4 8-A100 days, while FlowTok-H increases the budget slightly to 26.1 8-A100 days. In contrast, the most efficient text-as-condition model, PixArt- [chen2024pixartalpha], still demands 94.1 8-A100 days. Compared to CrossFlow [liu2024flowing], which also treats text as source distributions and requires 78.8 8-A100 days, FlowTok is much more efficient.
@@
Image-to-Text Generation. We evaluate image-to-text generation on COCO [lin2014microsoft] using the Karpathy split [karpathy2015deep], with results summarized in Tab. 3. To ensure a fair comparison, we categorize methods into two groups: direct flow from image to text distributions, which represents a new paradigm leveraging flow matching for direct image-to-text transformation, and other methods, considering only those trained without CIDEr optimization. Within the direct flow category, FlowTok-XL consistently outperforms its counterpart, CrossFlow [liu2024flowing], across most metrics. Specifically, FlowTok-XL achieves a BLEU-4 (B@4) score of 37.1, surpassing CrossFlow by 0.7, and a CIDEr score of 117.0, exceeding CrossFlow by 0.8. Moreover, compared to state-of-the-art methods from other paradigms, FlowTok-XL demonstrates competitive performance, highlighting direct flow matching as a promising approach for image-to-text generation. Notably, FlowTok performs image-to-text generation under the same formulation using compact 1D tokens, theoretically requiring fewer training resources and enabling faster sampling compared to paradigms that operate on 2D latents, as adopted by CrossFlow. However, a direct quantitative comparison is not possible, as CrossFlow has not released the corresponding checkpoint for evaluation.
@@ 5.3 Ablation Studies
Text Alignment Target. We first investigate the choice of alignment target for the projected text tokens in Tab. LABEL:tab:ablation_loss_target. A straightforward baseline is to directly apply average pooling (row 1) to the original CLIP text embedding along the channel dimension, reducing the dimensionality from 768 to 16 to match . However, this approach performs significantly worse compared to using a simple MLP to learn the alignment target (row 2), as inspired by prior works [grill2020bootstrap, oquab2023dinov2, yu2025representation]. We attribute this performance gap to the fact that adjacent channels in the CLIP text embedding are not necessarily correlated, and simple average pooling discards too much semantic information. In contrast, a learnable MLP mitigates this information loss, making it a more effective choice for defining the text alignment target.
Text Alignment Loss Function. Next, we examine the text alignment loss function in Tab. LABEL:tab:ablation_loss_function. Besides the contrastive loss adopted from [radford2021learning], we explore using a cosine similarity loss, similar to [yang2024depth]. Specifically, we compute the cosine similarity between the text tokens and the alignment target, applying a penalty to pairs with similarity below a threshold. Our experiments show that while both loss functions are effective, the contrastive loss achieves better performance.